If the discount yield on a 30-day Treasury bill is 4%, what is its bond-equivalent yield (investment rate)?
Answer #1

The below expression can be used to calculate the bond-equivalent yield:

Bond\, equivalent\, yield=\frac{365\times Discount\, yield}{360-\left ( Days\, to\, maturity\times Discount\, yield \right )}

The bond equivalent yield is calculated below:

Bond\, equivalent\, yield=\frac{365\times 0.04}{360-\left ( 30\times 0.04 \right )}

Bond\, equivalent\, yield=\frac{14.6}{358.8}

Bond\, equivalent\, yield=0.0407

Bond equivalent yield is 4.07%.